6. $7,500.00 deposited in an account earning 10.26%, for 8 years. Calculate the future value: 16,383.44 
...

6. $7,500.00 deposited in an account earning 10.26%, for 8 years. Calculate the future value:

  1. 16,383.44 

  2. 16,625.25 

  3. 16,935.25 

  4. 17,028.60


7. $7,500.00 deposited in an account earning 10.26%, for 8 years. Calculate the interest factor:

a. 2.4236

b. 2.1845

c. 2.0325

d. 1.9275

9. 7,500.00 deposited in an account earning 10.26% for 8 years, compounded weekly. Calculate the future value

a. 16,383.44

b. 16,625.25

c. 16,935.25

d. 17,028.60
Answer #1

6.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=$7500*(1.1026)^8

=$16,383.44(Approx).

7.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence interest factor=(1.1026)^8

=2.1845(Approx).

8.

We use the formula:
A=P(1+r/52)^52n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=7500*(1+0.1026/52)^(52*8)

=$7500*2.270480113

=$17028.60(Approx).